OpenStudy (anonymous):

A recipe that makes 3 dozen chocolate chip cookies calls for cups of flour. a) Exactly how much flour would you need for half the recipe? b) Exactly how much flour would you need to make 5 dozen cookies?

OpenStudy (anonymous):

Dude, you're posting all your homework questions. You're bound to fail the class.

OpenStudy (anonymous):

and how many cups of flour?

OpenStudy (anonymous):

I know! but I just don't understand fractions well. the other subjects aren't bad. I have to be able to explain them to a little kid and so i figure if I have the right answer I can try and work my way through it backwards. Sounds crazy but it makes sense to me :/

OpenStudy (anonymous):

What grade is this for?

OpenStudy (anonymous):

2 2/5 cups of flour. Its for college the course is called math for elementary teachers and my professor give us these take homes with word problems and we have to figure them out and then write it out so a little kid could understand this is the only one that has given me problems

OpenStudy (anonymous):

or nm, let's just leave it as a fraction

OpenStudy (anonymous):

Okay how to make this fraction easier to work with:

OpenStudy (anonymous):

Is that 2 cups + 2/5? If so, then you have 2*(5/5) cups + 2/5, which is 10/5 + 2/5. Since the denominators are the same, you can add these up, which gives you 12/5. When you're asked to find half that value, you can multiply the fraction by 1/2. 12/5 * 1/2 = 12 / 10.

OpenStudy (anonymous):

yea I do but she doesn't want us to I was thinking I had to turn it into an impropper fraction but im not sure

OpenStudy (anonymous):

Multiply you whole number (the 2) by the denominator (the 5) and add it to the numerator (the 2). (2*5)+2 = 12 This will be your numerator and your denominator will be 5. 12/5 cups

OpenStudy (anonymous):

o snap making some sense here :)

OpenStudy (anonymous):

As for the second part, you know 12/5 cups will make 3 dozens. You can divide that answer by 3 to find out how many cups you need for 1 dozen. 12/5 * 1/3 = 12/15 cups. Now, multiply that answer by 5 to find out how many you need for 5 dozens. 12/15 * 5 = 60/15, which is equal to 4.

OpenStudy (anonymous):

Now multiply 12/5 by 1/2. 12/5*(1/2) = 12/10 cups for 1.5 dozen cookies

OpenStudy (anonymous):

thank you both so very very much!

OpenStudy (anonymous):

plz read through this http://www.slideshare.net/sondrateer/fractions-1196964

OpenStudy (anonymous):

And look up mixed numbers as well for extra help.

OpenStudy (anonymous):

sounds good it'll help me for the test!